OpenStudy (anonymous):

A rectangle has an are of 24.5 square inches. The width of the rectangle is x + 4 *2 and the length is x. Use the quadratic formula to find the dimensions of the rectangle.

OpenStudy (anonymous):

A=l*w 24.5=(x + 4 *2)(x) 24.5=(x + 8)(x) 24.5=x^2 + 8x x^2 + 8x-24.5=0 Use the quadratic formula to find the dimensions : [-b∓√(b^2-(4*a*c)]/2a a=1 b=8 c=-24.5

OpenStudy (radar):

w=x+4*2 Given length l=x Given Area = 24.5 sq in. Given Area=lenght times width A=x(x +4*2)=x^2+8x=24.5 sq in. \[x ^{2}+8x=24.5\] \[x ^{2}+8x-24.5=0\]

OpenStudy (radar):

The equation is in the quadratic form and now can be solved using the quadratic formula